A compound formed by two Elements M and N the element N forms ccp and atoms of element M occupy 1/3 of tetrahedral voids. What is the formula of the compound? 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
In CCP crystalNumber of N atoms (Z) = 4Number of tetrahedral voids = 2Z = 8Number of M atoms = 1/3rd of 8 = 8/3N:M = 4: 8/3Formula, = N12M8orN3M2